The information below is to be used for questions 1-3 Month Labor Cost Employee Hours January 14,000 720 February 16,280 1,100 March 19,798 1,260 April 19,574 1,220 May 16,980 960 June 14,900 700 July 18,980 740 August 15,062 620

What is the fixed cost of labor?
Answer #1

Variable cost per unit = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (19,798-15,062)/(1,260-620)

= 4,736/640

= $7.4 per employee hour

Fixed cost = Highest activity cost - Highest activity x Variable cost per employee hour

= 19,798-1,260 x 7.4

= 19,798-9,324

= $10,474

the fixed cost of labor = $10,474

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!